OpenStudy (anonymous):

Determine the equations of any vertical asymptotes and the values of x for any holes in the graph of the following rational function. F(g)=(g-7/g^2-11g+28)

OpenStudy (anonymous):

x=7 is a hole. The other zero of the denominator is at x=4

OpenStudy (anonymous):

how did you figure that out ?

OpenStudy (anonymous):

the denom. is factorable: two roots x=7 and x=4 x=7 makes the function = 0/0

OpenStudy (anonymous):

zero in the denom. means a vertical asymptote, unless the numerator is zero also, then it's a hole.

OpenStudy (anonymous):

can you show how you figured it ?

OpenStudy (anonymous):

factor the denominator.

OpenStudy (anonymous):

(x- _?_) (x-_?_)

OpenStudy (anonymous):

where did you get 4?

OpenStudy (anonymous):

11-7?

OpenStudy (anonymous):

Did you factor the denominator?

OpenStudy (anonymous):

i dont know how thats why i need a drawing.

OpenStudy (anonymous):

aww c'mon, you've been factoring quadratics for years now, or at least several semesters... give it a shot..

OpenStudy (anonymous):

i dropped out. I have no hands on help.

OpenStudy (anonymous):

(x-4)(x-7)

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so you wouldnt do anything with G^2 or the g's in the equation?

OpenStudy (anonymous):

eg x^2 + 3x -18 factors (x+6)(x-3) 3*-6 = -18 6 -3 =3

OpenStudy (anonymous):